Suppose an investment offers to quadruple your money in 24 months (don’t believe it). What rate of return per quarter are you being offered? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value($4x)
P=present value($x say)

r=rate of interest
n=time period.(2*[12/3]=8)

4x=x(1+r/100)^(8)

4^(1/8)=(1+r/100)

(1+r/100)=1.1892

r=(1.1892-1)*100

=18.92%(Approx).